OpenStudy (anonymous):

A shop sells compact disks for $18 per piece. If each disk costs $10 to the shop owner how much profit percent does the shop owner earn?

OpenStudy (anonymous):

profit=8(18-10) profit%=8/cp*100=8/10*100=80%

OpenStudy (anonymous):

profit%=profit/costpriceX100 therefor profit=8/10x100=80%